# Coexistence of Topological Anderson Insulator and Multifractal Critical Phase in a Non-Hermitian Quasicrystal

### v1

@@ Abstract
Extended-critical-localized phase transitions-Quasiperiodic nonreciprocity also drives Anderson-type localization transitions. As shown in Figs. 1(a) and 1(b), the spectrum separates into three distinct regimes characterized by the fractal dimension $FD$ of the eigenstates. For $v=1.2$, all states are extended when $|\lambda|<1.2$, become multifractal critical in the intermediate region $1.2<|\lambda|<2$, and are fully localized for $|\lambda|>2$. Representative wave-function profiles are displayed in Figs. 1(d) and 1(e), where the critical states exhibit clear multifractal spatial structures.
To quantify the transition, we compute the mean inverse participation ratio and mean fractal dimension, $\mathrm{MIPR}=\sum_{n}(\mathrm{IPR}_{n})/L$ and $\mathrm{MFD}=\sum_{n}(\mathrm{FD}_{n})/L$. Figure 3 shows their dependence on $\lambda$ for different system sizes. In the extended regime, $\mathrm{MIPR}\rightarrow 0$ and $\mathrm{MFD}\rightarrow 1$ as $L$ increases, while in the localized regime $\mathrm{MIPR}$ remains finite and $\mathrm{MFD}\rightarrow 0$. In contrast, both quantities remain take intermediate values in the critical regime, consistent with multifractal scaling.
The localization transition can be determined analytically using Avila’s global theory for quasiperiodic $SL(2,\mathcal{C})$ cocycles [89]. The transfer matrix defines an analytic cocycle $(\alpha,T_{E}(\theta))$ and the Lyapunov exponent admits the phase-averaged representation
$\gamma(E)=\lim_{N\rightarrow\infty}\frac{1}{N}\int_{0}^{2\pi}\ln\left\|T_{E}\left[\theta+(N-1)\alpha\right]\cdots T_{E}(\theta)\right\|\frac{d\theta}{2\pi}.$ (12)
Analytically continuing $\theta\rightarrow\theta+iy$ with $y\in\mathcal{R}$, we obtain the exact result [82]
$\gamma(E)=\text{max}\left\{0,\ln\frac{|\lambda|}{2|w|}\right\}.$ (13)
Importantly, $\gamma(E)$ is independent of energy, implying the absence of mobility edges. For $|\lambda|>2|w|$, $\gamma(E)>0$ and all eigenstates are localized. For $|\lambda|<2|w|$, $\gamma(E)=0$, and the states are either extended or critical.
The distinction between extended and critical phases is controlled by the competition between $v$ and $\lambda$. When $|\lambda|<|v|$, the spectrum is absolutely continuous and all states are extended. For $|v|<|\lambda|<2|w|$, the spectrum becomes singular continuous [90], and all eigenstates are multifractal. Physically, the critical regime originates from incommensurate zeros in the effective hopping amplitudes when $|\lambda|>|v|$ [91, 92].
The system therefore exhibits three different phases
$\left\{\begin{array}[]{ll}\text{extended phase:}\quad&0<|\lambda|<|v|,\\ \text{critical phase:}\quad&|v|<|\lambda|<2|w|,\\ \text{localized phase:}\quad&|\lambda|>2|w|.\end{array}\right$. (14)
The resulting phase diagram is shown in Fig. 2(b). Comparing with the TAI diagram in Fig. 2(a), we uncover a finite parameter region where the topologically nontrivial phase overlaps with the multifractal critical regime. Along the representative cut $v=1.2$ [black dashed line in Fig. 2(b)], the system evolves as follows: starting from a trivial extended phase at $\lambda=0$, increasing $\lambda$ first induces a TAI phase with extended bulk states; further increasing $\lambda$ drives the bulk into a multifractal critical phase while the zero-energy edge modes persist; finally, for $|\lambda|>2|w|$, the system becomes localized and topologically trivial. Notably, the extended–critical phase transition at $|\lambda|=|v|$ coincides exactly with the real–complex spectral transition, establishing a direct correspondence between spectral topology and wave-function criticality. Moreover, the critical window shrinks as $|v|$ increases and disappears entirely for $|v|>2|w|$.
@@
Figure 3: (Color online) (a) Mean inverse participation ratio (MIPR) and (b) mean fractal dimension (MFD) as functions of the nonreciprocity strength $\lambda$ for different system sizes. The black dot-dashed and blue dashed lines mark the analytical phase boundaries separating the extended–critical and critical–localized regimes, respectively. Parameters: $v=1.2$ and $w=1$.
@@
Experimental realization.-Topolectrical circuits have emerged as a versatile platform for simulating topological band structures and non-Hermitian phenomena [13, 14, 15]. We propose a circuit implementation of the non-Hermitian quasicrystal, schematically shown in Fig. 4. In the circuit network, nodes represent lattice sites of the SSH chain. The capacitors $C_{1}$ and $C_{2}$ realize the intra- and intercell hopping amplitudes, respectively. Nonreciprocal intracell hopping is implemented using negative-impedance converters with current inversion (INICs), whose direction-dependent impedance $Z_{j}$ generates the asymmetric hopping terms. By appropriately designing $Z_{j}$s and $Z_{j}^{\prime}$, the circuit Laplacian $\bm{J}$ defined through $\bm{I=JV}$ (with $\bm{I}$ and $\bm{V}$ the nodal current and voltage vectors), reproduces the effective Hamiltonian in Eq. (1[82]. Topological and localization properties can be probed via impedance measurements. In particular, zero-energy edge modes manifest as pronounced peaks in the two-point impedance at the boundaries, while the admittance spectrum directly reveals the real–complex spectral transition and the localization regimes predicted in this work.
@@
Summary-We have introduced and analytically characterized a non-Hermitian SSH quasicrystal with quasiperiodically modulated nonreciprocity. The interplay between non-Hermiticity and quasiperiodicity gives rise to two intertwined phenomena: a disorder-induced non-Hermitian TAI and a cascade of extended–critical–localized transitions. Most notably, we uncover a finite parameter regime where the TAI phase coexists with a multifractal critical bulk—an effect absent in previously studied disordered topological systems. We establish exact analytical phase boundaries and demonstrate that the extended–critical transition coincides precisely with the real–complex spectral transition, revealing a direct correspondence between spectral feature and wave-function criticality. The proposed topolectrical circuit implementation offers an experimentally accessible route to probe these multifractal topological states. More broadly, quasiperiodic nonreciprocity provides a general mechanism for engineering hybrid quantum phases in which topology, non-Hermiticity, and criticality are intrinsically intertwined.

### v2

@@ IV Extended-critical-localized phase transitions
Quasiperiodic nonreciprocity also drives Anderson-type localization transitions. As shown in Figs. 1(a) and 1(b), the spectrum separates into three distinct regimes characterized by the fractal dimension $\Gamma$ of the eigenstates. For $v=1.2$, all states are extended when $|\lambda|<1.2$, become multifractal critical in the intermediate region $1.2<|\lambda|<2$, and are fully localized for $|\lambda|>2$. Representative wave-function profiles are displayed in Figs. 1(d) and 1(e), where the critical states exhibit clear multifractal spatial structures.
To quantify the transition, we compute the average values of the inverse participation ratio and fractal dimension, which are respectively defined as $\langle IPR\rangle=\sum_{n}\left[\mathrm{IPR}(\psi_{nR})\right]/L$ and $\langle D_{2}\rangle=\sum_{n}\left[D_{2}(\psi_{nR})\right]/L$. Figure 3(a) and 3(b) show their dependence on $\lambda$ for different system sizes under periodic boundary conditions, such that the influences of topological edge modes are excluded. In the extended regime, $\langle IPR\rangle\rightarrow 0$ and $\langle D_{2}\rangle\rightarrow 1$ as $L$ increases, while in the localized regime $\langle IPR\rangle$ remains finite and $\langle D_{2}\rangle\rightarrow 0$. In contrast, both quantities remain taking intermediate values in the critical regime, and the $\langle D_{2}\rangle$ behaves independently with the system size, consistent with multifractal scaling. Figure 3(c) further plots the $D_{2}$ values of all eigenstates, where the color indicates the fractal dimensions of eigenstates. We can see that there is no mobility edges in the spectrum, and the boundaries indicated by the color scale at $\lambda=1.2$ and $2$ are quite clear, consistent with the results presented in Fig. 3(a) and 3(b). For the multifractral critical phase, we further calculate the $D_{q}$ values of three representative eigenstates: $n=1$, 300300, and $610$, with $q$ varying from $2$ to $10$, as shown in Fig. 3(d). We can see that the $D_{q}$ values are $q$ dependent, but always remain a finite value satisfying $0<D_{q}<1$, as expected for the multifractal states.
@@
Figure 3: (Color online) (a) Average value of the inverse participation ratio $\langle IPR\rangle$ and (b) average value of the fractal dimension $\langle D_{2}\rangle$ as functions of the nonreciprocity strength $\lambda$ for different system sizes under periodic boundary conditions. The black dot-dashed and blue dashed lines mark the analytical phase boundaries separating the extended–critical and critical–localized regimes, respectively. (c) shows the $D_{2}$ values of all the eigenstates of the system with $N=610$, which are indicated by the color scale. (d) Fractal dimensions $D_{q}$ for the $1st$, 300300th, and 610610th eigenstate at $\lambda=1.5$ in (c) as a function of $q$. Parameters: $v=1.2$ and $w=1$.
The localization transition can be determined analytically by using Avila’s global theory for quasiperiodic cocycles [100, 101]. For the case with $|\lambda|<|v|$, the transfer matrix defines an analytic cocycle $(\alpha,T_{E}(\theta))$. However, if $|\lambda|\geq|v|$, the transfer matrix in Eq. (6) will become sigular due to the $(v+\lambda\cos\theta_{j})$ term in the denominator and the cocycle $(\alpha,T_{E}(\theta))$ also becomes singluar. Since the Avila’s global theory is developed to one-frequency analytic cocycles [100], we need to extend it to deal with the singular case here. To do so, we define the following new transfer matrix
$\widetilde{T}_{E}(\theta_{j})=(v+\lambda\cos\theta_{j})T_{E}(\theta_{j}).$ (15)
Then the $N$-step transfer matrix is $\widetilde{\mathcal{T}}_{N}=\left[\prod_{j=1}^{N}(v+\lambda\cos\theta_{j})\right]\mathcal{T}_{N}$, with $\mathcal{T}_{N}=\prod_{j=1}^{N}T_{E}(\theta_{j})$, and we have
$\ln\left\|\widetilde{\mathcal{T}}_{N}\right\|=\sum_{j=1}^{N}\ln|v+\lambda\cos\theta_{j}|+\ln\left\|\mathcal{T}_{N}\right\|.$ (16)
According to the definition of Lyapunov exponent in Eq. (7), the above formula leads to
$\widetilde{\gamma}(E)=\gamma(E)+\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}\ln|v+\lambda\cos\theta_{j}|.$ (17)
Since $\widetilde{T}_{E}(\theta_{j})$ is analytic, the transfer matrix defines an analytic $SL(2,\mathcal{R})$ cocycle, which is denoted as $(\alpha,\widetilde{T}_{E}(\theta_{j}))$. Then we can use the Avila’s theory to compute the $\widetilde{\gamma}(E)$. Due to the ergodicity, the Lyapunov exponent admits a phase-averaged representation
$\widetilde{\gamma}(E)=\lim_{N\rightarrow\infty}\frac{1}{2\pi N}\int_{0}^{2\pi}\ln\left\|\widetilde{T}_{E}(\theta_{j})\right\|d\theta.$ (18)
By replacing $\theta$ with $\theta+iy$ with $y\in\mathcal{R}$, we get the complexified Lyapunov exponent and have
$\gamma(E,y)=\widetilde{\gamma}(E,y)-\int_{0}^{2\pi}\ln|v+\lambda\cos(\theta+iy)|\frac{d\theta}{2\pi}.$ (19)
Setting $I(y)=\int_{0}^{2\pi}\ln|v+\lambda\cos(\theta+iy)|\frac{d\theta}{2\pi}$, one has
$I(y)=\left\{\begin{array}[]{ll}\ln\frac{\left|v\right|+\sqrt{v^{2}-\lambda^{2}}}{2},\qquad&\left|\lambda\right|<\left|v\right|;\\ |y|+\ln\frac{|\lambda|}{2},\qquad&\left|\lambda\right|\geq\left|v\right|.\end{array}\right$. (20)
Now since $\widetilde{T}_{E}(\theta_{j})$ is analytic, Avila’s global theory ensures that $\widetilde{\gamma}(E,y)$ is a continuous, convex, piecewise linear function with respect to $y$. For large $|y|$, the term $\lambda\cos(\theta+iy)$ behaves as
$\lambda\cos(\theta+iy)\sim\frac{\lambda}{2}e^{|y|}e^{\pm i\theta}.$ (21)
In this limit, the transfer matrix becomes asymptotically triangular. Consequently, for the Lyapunov exponent $\widetilde{\gamma}(E,y)$, we have
$\widetilde{\gamma}(E,y)=2|y|+\ln\left|\frac{\lambda^{2}}{4w}\right|+o(1),$ (22)
where $o(1)\rightarrow 0$ in the large $|y|$ limit. This leads to
$\gamma(E,y)=2|y|+\ln\left|\frac{\lambda^{2}}{4w}\right|-I(y),$ (23)
and we have
$\gamma(E,y)=\left\{\begin{array}[]{ll}2|y|+\ln\frac{\left|v\right|-\sqrt{v^{2}-\lambda^{2}}}{2\left|w\right|},\qquad&\left|\lambda\right|<\left|v\right|;\\ |y|+\ln\left|\frac{\lambda}{2w}\right|,\qquad&\left|\lambda\right|\geq\left|v\right|.\end{array}\right$. (24)
The physical Lyapunov exponent corresponds to $\gamma(E,0)$.
From Avila’s global theory, the physical Lyapunov exponent is obtained by matching the asymptotic behavior at large $|y|$, which gives
$\gamma(E)=\left\{\begin{array}[]{ll}\text{max}\left\{0,\ln\frac{\left|v\right|-\sqrt{v^{2}-\lambda^{2}}}{2\left|w\right|}\right\},\qquad&\left|\lambda\right|<\left|v\right|;\\ \text{max}\left\{0,\ln\left|\frac{\lambda}{2w}\right|\right\},\qquad&\left|\lambda\right|\geq\left|v\right|.\end{array}\right$. (25)
Remarkably, $\gamma(E)$ is independent of $E$. Therefore, all bulk states share the same localization behavior, and no mobility edge appears in the spectrum. If $\gamma(E)=0$, the system is in the extended or critical phase; while if $\gamma(E)>0$, the system is in the localized phase. The phase boundaries can be determined from the above formula by setting $\gamma(E)=0$, which leads to two critical lines
$|\lambda_{c1}|=2|w|\sqrt{\left|\frac{v}{w}\right|-1},\quad\text{and}\quad|\lambda_{c2}|=2|w|,$ (26)
which coincides at $|v|=2|w|$.
Note that if $|\lambda|<|v|<2|w|$, $\ln\frac{\left|v\right|-\sqrt{v^{2}-\lambda^{2}}}{2\left|w\right|}<0$, then according to Eq. (25), $\gamma(E)=0$ will always hold, implying that the system will in the extended or critical phase in this regime. On the other hand, if $|v|\leq|\lambda|<2|w|$, $\gamma(E)$ is determined by $\ln\left|\frac{\lambda}{2w}\right|$, then the critical disorder strength is $|\lambda_{c2}|=2|w|$. Thus, when $|\lambda|<2|w|$, $\gamma(E)=0$, the system is in the extended or critical phase. If $|\lambda|>2|w|$, $\gamma(E)>0$, and all eigenstates are localized. The distinction between extended and critical phases is controlled by the competition between $v$ and $\lambda$. When $|\lambda|<|v|$, the spectrum is absolutely continuous and all states are extended. For $|v|<|\lambda|<2|w|$, the spectrum becomes singular continuous [102], and all eigenstates are multifractal. Physically, the critical regime originates from incommensurate zeros in the effective hopping amplitudes when $|\lambda|>|v|$ [103, 104]. So, for the region with $|v|<2|w|$, the system therefore exhibits three different phases
$\left\{\begin{array}[]{lll}\text{extended phase:}\quad&0<|\lambda|<|v|,\\ \text{critical phase:}\quad&|v|<|\lambda|<2|w|,\\ \text{localized phase:}\quad&|\lambda|>2|w|.\end{array}\right$. (27)
The resulting phase diagram is shown in Fig. 2(b). Comparing with the topological phase diagram in Fig. 2(a), we uncover a finite parameter region where the topologically nontrivial phase overlaps with the multifractal critical regime. Along the representative cut $v=1.2$, as indicated by the black dashed line in Fig. 2(b), the system evolves as follows: starting from a trivial extended phase at $\lambda=0$, increasing $\lambda$ first induces a TAI phase with extended bulk states; further increasing $\lambda$ drives the bulk into a multifractal critical phase while the zero-energy edge modes persist; finally, for $|\lambda|>2|w|$, the system becomes localized and topologically trivial. Notably, the extended–critical phase transition at $|\lambda|=|v|$ coincides exactly with the real–complex spectral transition, establishing a direct correspondence between the spectral property and wave-function criticality. Moreover, the critical window shrinks as $|v|$ increases and disappears entirely when $|v|>2|w|$.
For the regime $|v|>2|w|$, the system exhibits the following two phases
$\left\{\begin{array}[]{ll}\text{extended phase:}\quad&0<|\lambda|<2|w|\sqrt{\left|\frac{v}{w}\right|-1},\\ \text{localized phase:}\quad&|\lambda|>2|w|\sqrt{\left|\frac{v}{w}\right|-1},\end{array}\right$. (28)
which are separated by $|\lambda_{c1}|$. This phase boundary is plotted in Fig. 2(b), as indicated by the black solid line separating the blue and red region.
Above we have analytically determined the phase boundaries for the extended-critical-localized phases in the SSH model nonreciprocal quasiperiodic intracell hopping by extending the Avila’s theory, the results are in good agreement with the numeric calculations. We also reveal the coexistence of multifractal critical phase and the nontrivial TAI phase induced by the nonreciprocal quasiperiodic modulation in the model, which have not been reported before.
V Experimental scheme in electric circuits
@@
Topolectrical circuits have emerged as a versatile platform for simulating topological band structures and non-Hermitian phenomena [13, 14, 15]. We propose a circuit implementation of the non-Hermitian quasicrystal, schematically shown in Fig. 4. In the circuit network, nodes represent lattice sites of the SSH chain. The capacitors $C_{1}$ and $C_{2}$ realize the intra- and intercell hopping amplitudes, respectively. Nonreciprocal intracell hopping is implemented using negative-impedance converters with current inversion (INICs), whose direction-dependent impedance $Z_{j}$ generates the asymmetric hopping terms. By appropriately designing $Z_{j}$s and $Z_{j}^{\prime}$, the circuit Laplacian $\bm{J}$ defined through $\bm{I=JV}$ (with $\bm{I}$ and $\bm{V}$ the nodal current and voltage vectors), reproduces the effective Hamiltonian in Eq. (1)(see Appendix). Topological and localization properties can be probed via impedance measurements. In particular, zero-energy edge modes manifest as pronounced peaks in the two-point impedance at the boundaries, while the admittance spectrum directly reveals the real–complex spectral transition and the localization regimes predicted in this work.
@@ VI Summary
We have introduced and analytically characterized a non-Hermitian SSH quasicrystal with quasiperiodically modulated nonreciprocity. The interplay between non-Hermiticity and quasiperiodicity gives rise to two intertwined phenomena: a disorder-induced non-Hermitian TAI and a cascade of extended–critical–localized transitions. Most notably, we uncover a finite parameter regime where the TAI phase coexists with a multifractal critical bulk—an effect absent in previously studied disordered topological systems. We establish exact analytical phase boundaries and demonstrate that the extended–critical transition coincides precisely with the real–complex spectral transition, revealing a direct correspondence between spectral feature and wave-function criticality. The proposed topolectrical circuit implementation offers an experimentally accessible route to probe these multifractal topological states. More broadly, quasiperiodic nonreciprocity provides a general mechanism for engineering hybrid quantum phases in which topology, non-Hermiticity, and criticality are intrinsically intertwined.
Appendix
VI.1 The condition for a real spectrum
Under open boundary conditions, the model Hamiltonian in Eq. (1) can be written as a tridiagonal matrix describing a one-dimensional nonreciprocal lattice. Defining the nearest-neighbor hopping amplitudes
$t_{2j-1}=v+\lambda\cos\theta_{j},\quad t_{2j-1}^{\prime}=v-\lambda\cos\theta_{j},\quad t_{2j}=t_{2j}^{\prime}=w,$ (A1)
where $t_{n}$ ($t^{\prime}_{n}$) is the backward (forward) hopping amplitude between the $n$th and the $(n+1)$th lattice sites. he Hamiltonian takes the form
$H=\left(\begin{array}[]{cccccc}0&t_{1}&0&0&\cdots&0\\ t^{\prime}_{1}&0&t_{2}&0&\cdots&0\\ 0&t^{\prime}_{2}&0&t_{3}&\cdots&0\\ 0&0&t_{3}^{\prime}&0&t_{4}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&\cdots&0\end{array}\right)=\left(\begin{array}[]{cccccc}0&v+\lambda\cos\theta_{1}&0&0&\cdots&0\\ v-\lambda\cos\theta_{1}&0&w&0&\cdots&0\\ 0&w&0&v+\lambda\cos\theta_{2}&\cdots&0\\ 0&0&v-\lambda\cos\theta_{2}&0&w&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&\cdots&0\end{array}\right).$ (A2)
All the elements in the matrix are real and we have $t_{j}\neq t^{\prime}_{j}$, implying that the matrix is non-Hermitian.
Assume that $t_{j}t^{\prime}_{j}>0$, we construct a diagonal matrix
$D=\left(\begin{array}[]{cccccc}d_{1}&0&0&0&\cdots&0\\ 0&d_{2}&0&0&\cdots&0\\ 0&0&d_{3}&0&\cdots&0\\ 0&0&0&d_{4}&0&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&\cdots&d_{L}\\ \end{array}\right),$ (A3)
with
$d_{n}=\left\{\begin{array}[]{c}1,\qquad j=1\\ \sqrt{\frac{t^{\prime}_{n-1}t^{\prime}_{n-2}\cdots t^{\prime}_{1}}{t_{n-1}t_{n-2}\cdots t_{1}}},\qquad n=2,3,\cdots,L\end{array}\right$. (A4)
Under the similarity transformation $h=D^{-1}HD$, the Hamiltonian becomes
$h=\left(\begin{array}[]{cccccc}0&sgn(t_{1})\sqrt{t_{1}t^{\prime}_{1}}&0&0&\cdots&0\\ sgn(t_{1})\sqrt{t_{1}t^{\prime}_{1}}&0&sgn(t_{2})\sqrt{t_{2}t^{\prime}_{2}}&0&\cdots&0\\ 0&sgn(t_{2})\sqrt{t_{2}t^{\prime}_{2}}&0&sgn(t_{3})\sqrt{t_{3}t^{\prime}_{3}}&\cdots&0\\ 0&0&sgn(t_{3})\sqrt{t_{3}t^{\prime}_{3}}&0&sgn(t_{4})\sqrt{t_{4}t^{\prime}_{4}}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&\cdots&0\\ \end{array}\right).$ (A5)
Because condition $t_{j}t^{\prime}_{j}>0$ guarantees, all square roots are real and $h$ is Hermitian. Since similarity transformations preserve the spectrum, $H$ and $h$ share identical eigenvalues. Therefore, the spectrum of $H$ is entirely real whenever the condition $t_{j}t^{\prime}_{j}>0$ holds.
For the present model,
$t_{2j-1}t_{2j-1}^{\prime}=(v+\lambda\cos\theta_{j})(v-\lambda\cos\theta_{j})=v^{2}-\lambda^{2}\cos^{2}\theta_{j}.$ (A6)
The condition $t_{j}t^{\prime}_{j}>0$ for all $j$ requires
$v^{2}-\lambda^{2}\cos^{2}\theta_{j}>0\quad\forall j,$ (A7)
which is satisfied if and only if
$|\lambda|<|v|.$ (A8)
Thus, the critical point for the real-to-complex spectral transition is
$|\lambda|=|v|.$ (A9)
For $|\lambda|>|v|$, the similarity transformation breaks down because some $t_{n}$ and $t_{n}^{\prime}$ become negative, and complex eigenvalues necessarily emerge.
The above similarity transformation implies that the Hamiltonian is pseudo-Hermitian. Indeed,
$D^{2}H^{\dagger}D^{-2}=D^{2}(DhD^{-1})^{\dagger}D^{-2}=DhD^{-1}=H,$ (A10)
which yields $H^{\dagger}=\eta^{-1}H\eta$ with $\eta=D^{2}$. Thus, the Hamiltonian is pseudo-Hermitian. Therefore, $H$ is pseudo-Hermitian and admits a real spectrum in the parameter regime $|\lambda|<|v|$.
Notably, this argument does not depend on the specific functional form of $t_{n}$ and applies generally to one-dimensional nonreciprocal tridiagonal Hamiltonians [93].
VI.2 Experimental scheme using topolectrical circuits
The non-Hermitian quasicrystal proposed in the main text can be experimentally realized using topolectrical circuits, which have emerged as a versatile platform for simulating non-Hermitian band structures and topological phases. The circuit configuration is shown in Fig. 4 of the main text. The numbered black dots denote circuit nodes corresponding to lattice sites, where external voltage sources or impedance measurements can be applied.
The circuit consists of capacitors and negative-impedance converters with current inversion (INICs), which generate direction-dependent couplings and thereby emulate nonreciprocal hopping. The INIC structure (lower panel of Fig. 4) comprises two resistors, one capacitor, and an operational amplifier. Importantly, its effective impedance depends on the direction of current flow: it behaves as a negative capacitance for one orientation and as a positive capacitance for the opposite orientation, thus realizing asymmetric hopping amplitudes.
For a node $j$ in the circuit, let $I_{j}$ and $V_{j}$ denote the input current and voltage, respectively. Kirchhoff’s current law gives
$I_{j}=\sum_{i}Y_{ji}(V_{j}-V_{i})+X_{j}V_{j},$ (A11)
where $Y_{ji}$ is the admittance between nodes $j$ and $i$, and $X_{j}$ denotes the shunt admittance at node $j$. Collecting all nodes, the voltage–current relation takes the matrix form
$\bm{I}=\bm{JV},$ (A12)
where $\bm{J}$ is the circuit Laplacian. The circuit Laplacian can be used to model the tight-binding lattices Hamiltonian.
For the electrical circuit shown in Fig. (4), the Laplacian is
$\bm{J}=i\omega\left(\begin{array}[]{cccccc}C_{1}+Z_{1}-C_{1}^{\prime}&-(C_{1}+Z_{1})&0&0&0&\cdots\\ -(C_{1}-Z_{1})&C_{1}+C_{2}-Z_{1}-C_{2}^{\prime}&-C_{2}&0&0&\cdots\\ 0&-C_{2}&C_{1}+C_{2}+Z_{2}-C_{3}^{\prime}&-(C_{1}+Z_{2})&0&\cdots\\ 0&0&-(C_{1}-Z_{2})&C_{1}+C_{2}-Z_{2}-C_{4}^{\prime}&-C_{2}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&\cdots&C_{1}-Z_{N}-C_{L}^{\prime}\end{array}\right).$ (A13)
This structure reproduces the tight-binding Hamiltonian under open boundary conditions, up to the overall factor $i\omega$.
To eliminate the diagonal (on-site) terms and obtain a purely off-diagonal SSH-type structure, we choose
$C_{1}^{\prime}=C_{1}+Z_{1},\qquad C_{L}^{\prime}=C_{1}-Z_{N},\qquad C_{2j}^{\prime}=C_{1}+C_{2}-Z_{j}\quad C_{2j-1}^{\prime}=C_{1}+C_{2}+Z_{j}\quad for\quad 1<j<N.$ (A14)
The Laplacian then reduces to
$\bm{J}=i\omega\left(\begin{array}[]{cccccc}0&-(C_{1}+Z_{1})&0&0&0&\cdots\\ -(C_{1}-Z_{1})&0&-C_{2}&0&0&\cdots\\ 0&-C_{2}&0&-(C_{1}+Z_{2})&0&\cdots\\ 0&0&-(C_{1}-Z_{2})&0&-C_{2}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&\cdots&0\end{array}\right).$ (A15)
Identifying
$C_{1}\leftrightarrow v,\qquad C_{2}\leftrightarrow w,\qquad Z_{j}=\lambda\cos(2\pi\alpha j+\phi),$ (A16)
the circuit Laplacian becomes directly proportional to the non-Hermitian quasicrystal Hamiltonian discussed in the main text.
The admittance spectrum of the circuit corresponds to the energy spectrum of the Hamiltonian. In particular, zero-admittance resonances signal topological edge modes. The energy spectrum is obtained from the admittance spectrum of the circuit and can be used to detect the spectral properties reported in this work.
Acknowledgements.